OpenStudy (anonymous):

Find the GCF: 18x^3 and 30x^5

OpenStudy (anonymous):

1.) First divide the numbers into primes: 18x^3 = 2 * 3 * 3 * x * x * x 30x^5 = 2 * 3 * 5 * x * x * x * x * x 2.) Circle what they have in common: 2 * 3 * x * x * x 3.) Multiply, Greatest Common Factor = 6x^3

OpenStudy (anonymous):

 GCF:6x^3